Domain and renge f(x)=1/2x-p

1
Expert's answer
2022-02-08T07:40:44-0500

Let us find the domain and range of "f(x)=\frac{1}{2x-p}."

It follows that the domain is equal to the set "\R\setminus\{\frac{p}2\}."

If "y=0" then the equation "y=\frac{1}{2x-p}" has no solution. If "y\ne 0" then the equation "y=\frac{1}{2x-p}" is equivalent to "2x-p=\frac{1}{y}", and hence has a solution "x=\frac{1}{2y}+\frac{p}2" . Therefore, "f(x)=f(\frac{1}{2y}+\frac{p}2)=y." We conclude that the range of the function "f" is equal to the set "\R\setminus\{0\}."